# Cohomological rank functions and syzygies of projective bundles on abelian varieties

### v3

@@ 1. Introduction
Let $X$ be an abelian variety over the complex number field $\mathbb{C}$. Given $(X,l)$ a polarized abelian variety and an object $\mathcal{F}\in D^{b}(X)$, Jiang–Pareschi [JP20] and Caucci [Cau20] define the cohomological rank functions associated to $\mathcal{F}$ and $l$
@@
$\beta(l)\coloneq\inf\{x\in\mathbb{{Q}}^{+}\>|\>h^{1}_{\mathcal{I}_{x},l}(x)=0\},$
@@
  • $\beta(l)<\frac{1}{2}$ if, and only if, for every $L$ with $c_{1}(L)=0$, we have that $L$ is very ample and normally generated (see [JP20, Cor. E]);
  • if $\beta(l)<\frac{1}{p+2}$, then, for every $L$ with $c_{1}(L)=0$, we have that $L$ is very ample and satisfies property $N_{p}$ (see [Cau20, Thm. 1.1]);
@@
Let $X$ be an abelian variety of dimension $g$ over the complex number $\mathbb{C}$. Let
$0=E_{0}\subset E_{1}\subset\cdots\subset E_{\mathfrak{r}}=E$
be a filtration of subspaces of vector bundles $E$ with $\mathrm{rk}(E_{i})=i$ on $X$ such that
[…]
Let $Y:=\mathbb{P}_{X}(E)$ be the associated projective vector bundle with a projective morphism $\pi:Y\rightarrow X$. We denote by $\mathcal{O}_{Y}(1)$ the tautological line bundle on $Y$. Let $p\geq 0$. Let $A$ be an ample line bundle on $X$ with
@@
Property $N_{p}^{0}$ is equivalent to property $N_{p}$. So, our first main theorem, Theorem A generalizes Chintapalli’s main result on property $N_{p}$ [Chi, Theorem 1.2]. As a corollary (see Corrollary 4.5), we get Chiantipalli original result that $\mathcal{O}_{Y}(1)\otimes\pi^{*}A^{\otimes n}$ satisfies $N_{p}$ for every $n\geq p+3$.
Projective normality and normal presentation on curves were initially studied by Castelnuova, and later by Mattuck, Fujita and St-Donat. Mark Green united these concepts as special cases of the $N_{p}$-properties, $p$-th syzygy property for $p\geq 0$, see [Gre84a, Gre84b, Gre89]. In particular, Green generalized the result by Castelnuovo, showing that a line bundle of degree at least $2g+1+p$ on a smooth curve of genus $g$ satisfies $N_{p}$-property. In case of higher dimensional varieties, Ein and Lazarsfeld [EL93], among many others, proved that adjoint linear series associated to a very ample line bundle on a smooth projective variety satisfies the property $N_{p}$. Park [Par07] generalized the result by Ein–Lazarsfeld to projective bundles over smooth projective varieties, under certain hypothesis.
As for abelian varieties, for an ample line bundle $A$ on an abelian variety, Lazarsfeld made a conjecture that $A^{n}$ satisfies $N_{p}$-properties if $n\geq p+3$ for $p\geq 0$. This conjecture was proved and generalized to property $N_{p}^{r}$ by Pareschi [Par00] in characteristic 00, where it is shown that $A^{n}$ satisfies $N_{p}^{r}$-property if $n\geq(p+r+3)/(r+1)$. Chintapalli and Iyer [CI14] proved the Pareschi result on property $N_{p}$ for hyperelliptic varieties, and Chintapalli [Chi, Theorem 1.2] extended the result to projective bundles over abelian varieties.
@@
Suppose $(Y,X,\pi)$ as before. Let $A$ be an ample line bundle on $X$ with no base divisor. Let $p\geq 1$. If $\mathcal{I}_{o}\left\langle\frac{r+1}{p+r+2}c_{1}(A)\right\rangle$ is M-regular, then $\mathcal{O}_{Y}(1)\otimes\pi^{*}A$ satisfies $N_{p}^{r}$-property.
@@
We utilize the work of Ito [Ito22] on the M-regularity of $\mathbb{Q}$-twisted sheaves to prove Theorem B. As a corollary of Theorem B, we have the following:
@@
Let $X$ and abelian varieties, $E$ a vector bundle as in the statement of Theorem B, and $A$ an ample line bundle on $X$. If $n\geq(p+r+2)/(r+1)$, then $\mathcal{O}_{Y}(1)\otimes\pi^{*}A^{\otimes n}$ satisfies $N_{p}^{r}$ for $r\geq 0$, $p\geq 1$.
Ito [Ito22] refines results of Pareschi–Popa [PP03, PP11] on global generation and surejectivity of multiplication maps of global sections of coherent sheaves on abelian varieties, using the theory of M-regularity of $\mathbb{Q}$-twisted sheaves. Furthermore, he shows the implication of the M-regularity of a suitable $\mathbb{Q}$-twisted sheaf to property $N_{p}$ and jet-ampleness for an ample line bundle on abelian varieties.
Let $X$ be a nodal curve, with one node, and $X_{0}$ its normalization. Desingularization $\widetilde{J}(X)$ of compactified Jacobian $\overline{J}(X)$ can be expressed as a $\mathbb{P}^{1}$-bundle $\mathbb{P}(E)$ over desingularization $\widetilde{J}(X_{0})$ of compactified Jacobian $J(X_{0})$, where $E$ is given by direct sum of degree zero line bundles on desingularization $\widetilde{J}(X_{0})$. In [Chi, Theorem 5.4], Chintapalli applied his result, Corrollary 4.5 to desingularization of compactified Jacobians by setting $X=J(X_{0})$ with the direct sum $E$ of degree zero line bundles.
In the last section, §5, we provide an application of our result to desingularization of compactified Jacobians, generalizing [Chi, Theorem 5.4].
@@
We work on a complex number field $\mathbb{C}$. We use standard notation for abelian varieties. If $X$ is an abelian variety, we denote by $\hat{X}$ its dual abelian variety. Given an element $\alpha\in\hat{X}$, we let $P_{\alpha}$ be the corresponding line bundle in $\operatorname{Pic}^{0}(X)$, that is $P_{\alpha}=\mathscr{P}_{|X\times\{\alpha\}}$, where $\mathscr{P}$ is the reduced Poincaré line bundle on $X\times\hat{X}$. Throughout this paper, we may use $\alpha$ to indicate $P_{\alpha}$. Given a integer $n$, we let $n_{X}:X\rightarrow X$ be the "multiplication by $n$" map. When $n$ is positive, we denote by $X[n]\coloneqq\operatorname{Ker}n_{X}$: the group-subscheme of $n$-torsion points of $X$. For every $x\in X$ we let $t_{x}:X\rightarrow X$ be the map defined by $a\mapsto a+x$. Given a polarization $l$ on $X$, we let $\varphi_{l}:X\rightarrow\hat{X}$ be the isogeny defined by $x\mapsto t_{x}^{*}L\otimes L^{-1}$, where $L$ is an ample line bundle with $c_{1}(L)=l$. It is a standard fact about abelian varieties that this does not depends from the choice of $L$. We will denote the kernel of $\varphi_{l}$ by $K(l)$. The identity element of $X$ is denoted simply by 0.
@@ 2.1. Syzygies of projective varieties
Let $Z$ be an algebraic variety over an algebraically closed field $\mathbb{K}$ and let $L$ be an ample invertible sheaf on $Z$. Let $S_{L}$ be the symmetric algebra over $H^{0}(Z,L)$. Then section ring associated to $L$,
@@
is a finitely generated graded $S_{L}$-algebra and as such it admits a minimal free resolution
(2.1) $E_{\bullet}=0\rightarrow\cdots\xrightarrow{f_{p+1}}E_{p}\xrightarrow{f_{p}}\cdots\xrightarrow{f_{2}}E_{1}\xrightarrow{f_{1}}E_{0}\xrightarrow{f_{0}}R_{L}\rightarrow 0$
with $E_{i}\simeq\bigoplus_{j}S_{L}(-a_{ij})$, $a_{ij}\in\mathbb{Z}_{>0}$. Following Green [Gre84a], we say that $L$ satisfies property $N_{p}$ for some nonnegative integer $p$, if, in the notations above,
$E_{0}=S_{L}$
@@ 2.2. $\mathbb{Q}$-twisted sheaves
$V^{i}\left(X,\mathcal{F}\left\langle\frac{a}{b}l\right\rangle,L\right)\coloneqq\{\alpha\in\hat{X}\>|\>h^{i}(X,b_{X}^{*}\mathcal{F}\otimes L^{ab}\otimes P_{\alpha})=0\}.$
If we change the representative of $l$ we obtain a new locus that is a translate of $V^{i}\left(X,\mathcal{F}\left\langle\frac{a}{b}l\right\rangle,L\right)$ by an element in $\hat{X}$. In particular, the dimension of these loci does not depend from the choice of the line bundle $L$ and we can introduce the following definition.
@@
We say that $\mathcal{F}\left\langle\frac{a}{b}l\right\rangle$ is GV if $\operatorname{codim}V^{i}(X,\mathcal{F}\left\langle\frac{a}{b}l\right\rangle,L)\geq i$ for every $i\geq 1$.
We say that $\mathcal{F}\left\langle\frac{a}{b}l\right\rangle$ is M-regular if $\operatorname{codim}V^{i}(X,\mathcal{F}\left\langle\frac{a}{b}l\right\rangle,L)\geq i+1$ for every $i\geq 1$.
We say that a $\mathbb{Q}$-twisted sheaf $\mathcal{F}\left\langle\frac{a}{b}l\right\rangle$ satisfies I.T.(0) if its cohomological support loci are empty for every $i>0$.
@@ 2.3. Cohomological rank functions on abelian varieties
We see that the value of $h^{i}_{\mathcal{F},l}(x)$ depends only of $\mathcal{F}\langle xl\rangle$. Furthermore, we have that a $\mathbb{Q}$-twisted sheaf $\mathcal{F}\langle xl\rangle$ satisfies I.T.(0) if, and only if, for every $i>0$, we have that $h^{i}{\mathcal{F},l}(x)=0$. In addition to this, cohomological rank functions have been used to define invariants associated to polarizations of abelian varieties. Let $p\in X$ a closed point, we denote by $\mathcal{I}_{p}$ its associated sheaf of ideals. Given $l$ a polarization on $X$, we have that the quantity
@@
(2.6) $k(l)=\frac{\beta(l)}{1-\beta(l)}.$
@@ 3. Ampleness of Tautological Bundles
Let $Z$ be a smooth projective complex variety. Let $V$ be a vector bundle of rank $\mathfrak{r}$ on $Z$. The following is the definition of $k$-jet ampleness [BDRS99, §1.2].
@@
A vector bundle $V$ is 00-jet ample if and only if it is globally generated. In addition, according to [BP08a, Theorem 4.2], if $V$ is a $1$-jet vector bundle on $Z$, then $V$ is very ample on $Z$.
Let $\mathbb{P}_{Z}(V)$ be the associated projective bundle with the projection morphism $\pi:\mathbb{P}_{Z}(V)\rightarrow Z$. Let $\mathcal{O}_{\mathbb{P}(V)}(1)$ be the tautological line bundle of $\mathbb{P}_{Z}(V)$.
Definition 3.2.
Thus, given a very ample vector bundle $V$ on $Z$, it make sense to give conditions ensuring property $N_{p}^{r}$ holds for the embedding induced by $\mathcal{O}_{\mathbb{P}(V)}(1)$. In [Par07] a cohomological criterion for property $N_{p}^{r}$ for the tautological bundle of a projective bundle $\mathbb{P}(V)\rightarrow Z$ is given. Before recalling the statement we need to fix some notation.
To this aim let $V$ be a very ample vector bundle on $Z$. In particular, we know that $Y$ is globally generated and we have a short exact sequence
$0\rightarrow M_{V}\longrightarrow H^{0}(Z,V)\longrightarrow V\rightarrow 0,$
where $M_{V}$ denotes the kernel of the evalutaion map $H^{0}(Z,V)\otimes\mathcal{O}_{V}\rightarrow V$.
@@
This is essentially [Par07, Prop. 3.2] where it is proven that the vanishing
@@
for every $l+m=i$ implies the vanishing
$H^{k}\left(Z,\bigwedge^{i}M_{V}\otimes V^{\otimes h}\right)=0.\qed$

### v4

@@ Abstract.
Key words and phrases:
Projective bundles, Syzygies, Abelian varieties
2020 Mathematics Subject Classification:
13D02, 14C20, 14F17
All authors contributed equally to this work.
@@ 1. Introduction
Let $X$ be an abelian variety over the field of complex numbers $\mathbb{C}$. Given $(X,l)$ a polarized abelian variety and an object $\mathcal{F}\in D^{b}(X)$, Jiang–Pareschi [JP20] and Caucci [Cau20] define the cohomological rank functions associated to $\mathcal{F}$ and $l$
@@
$\beta(l):=\inf\{x\in\mathbb{{Q}}^{+}\>|\>h^{1}_{\mathcal{I}_{x},l}(x)=0\},$
@@
  • $\beta(l)<\frac{1}{2}$ if, and only if, for every $L$ with $c_{1}(L)=l$, we have that $L$ is normally generated (see [JP20, Cor. E]);
  • if $\beta(l)<\frac{1}{p+2}$, then, for every $L$ with $c_{1}(L)=l$, we have that $L$ satisfies property $N_{p}$ (see [Cau20, Thm. 1.1]);
We recall that a vector bundle $E$ on an abelian variety $X$ is called homogeneous if, and only if, $t_{x}^{*}E\simeq E$ for all $x\in A$, where $t_{x}:X\rightarrow X$ denotes the "translation by $x$". This condition is known to be equivalent to $E$ having a filtration
$0=E_{0}\subset E_{1}\subset\cdots\subset E_{\mathfrak{r}}=E$
of subvector bundles with $\mathrm{rk}(E_{i})=i$ and
$E_{t}/E_{t-1}\in\mathrm{Pic}^{0}(X)\;\text{for}\;1\leq t\leq\mathfrak{r}.$
@@
Let $X$ be an abelian variety of dimension $g$ over the complex number $\mathbb{C}$ and consider an homogeneous vector bundle $E$ on $X$. Let $Y:=\mathbb{P}_{X}(E)$ be the associated projective vector bundle with a projective morphism $\pi:Y\rightarrow X$. We denote by $\mathcal{O}_{Y}(1)$ the tautological line bundle on $Y$. Let $p\geq 0$. Let $A$ be an ample line bundle on $X$ with
@@
Property $N_{p}^{0}$ is equivalent to property $N_{p}$. So, our first main theorem, Theorem A generalizes Chintapalli’s main result on property $N_{p}$ [Chi19, Theorem 1.2], however we were told by A. Ito that Theorem A above is a particular case of the main theorem in the appendix of [Ray24], where the filtration assumption is not needed. As a corollary (see Corollary 4.5), we get Chiantipalli original result that $\mathcal{O}_{Y}(1)\otimes\pi^{*}A^{\otimes n}$ satisfies $N_{p}$ for every $n\geq p+3$.
Projective normality and normal presentation on curves were initially studied by Castelnuovo, and later by Mattuck, Fujita and St-Donat. Mark Green merged these concepts as special cases of the $N_{p}$-properties, $p$-th syzygy property for $p\geq 0$, see [Gre84a, Gre84b, Gre89]. In particular, Green generalized the result by Castelnuovo, showing that a line bundle of degree at least $2g+1+p$ on a smooth curve of genus $g$ satisfies $N_{p}$-property. In case of higher dimensional varieties, Ein and Lazarsfeld [EL93], among many others, proved that adjoint linear series associated to a very ample line bundle on a smooth projective variety satisfies the property $N_{p}$. Park [Par07] generalized the result by Ein–Lazarsfeld to projective bundles over smooth projective varieties, under certain hypothesis.
As for abelian varieties, for an ample line bundle $A$ on an abelian variety, Lazarsfeld made a conjecture that $A^{n}$ satisfies $N_{p}$-properties if $n\geq p+3$ for $p\geq 0$. This conjecture was proved and generalized to property $N_{p}^{r}$ by Pareschi [Par00] in characteristic 00, where it is shown that $A^{n}$ satisfies $N_{p}^{r}$-property if $n\geq(p+r+3)/(r+1)$. Chintapalli and Iyer [CI14] proved the Pareschi result on property $N_{p}$ for hyperelliptic varieties, and Chintapalli [Chi19, Theorem 1.2] extended the result to homogeneous projective bundles over abelian varieties, that is to projectivizations of homogenous vector bundles.
@@
Suppose $(Y,X,\pi)$ as before. Let $A$ be an ample line bundle on $X$. Let $p\geq 1$. If $\mathcal{I}_{o}\left\langle\frac{r+1}{p+r+2}c_{1}(A)\right\rangle$ is M-regular, then $\mathcal{O}_{Y}(1)\otimes\pi^{*}A$ satisfies $N_{p}^{r}$-property.
@@
We utilize the work of Ito [Ito22] on the M-regularity of $\mathbb{Q}$-twisted sheaves to prove Theorem B.
Ito [Ito22] refines results of Pareschi–Popa [PP03, PP11] on global generation and surjectivity of multiplication maps of global sections of coherent sheaves on abelian varieties, using the theory of M-regularity of $\mathbb{Q}$-twisted sheaves. Furthermore, he shows the implication of the M-regularity of a suitable $\mathbb{Q}$-twisted sheaf to property $N_{p}$ and jet-ampleness for an ample line bundle on abelian varieties.
As a corollary of Theorem B, we have the following:
@@
Let $X$ an abelian variety, $E$ a vector bundle as in the statement of Theorem B, and $A$ an ample line bundle on $X$ with no base divisor. If $n\geq(p+r+2)/(r+1)$, then $\mathcal{O}_{Y}(1)\otimes\pi^{*}A^{\otimes n}$ satisfies $N_{p}^{r}$ for $r\geq 0$, $p\geq 1$.
Our Corollary C gives refined results of Chintapalli, since it implies that $\mathcal{O}_{Y}(1)\otimes\pi^{*}A^{\otimes n}$ satisfies $N_{p}$ for $n\geq p+2$.
Let $X$ be a nodal curve, with one node, and $X_{0}$ its normalization. Desingularizations $\widetilde{J}(X)$ of compactified Jacobian $\overline{J}(X)$ can be expressed as a $\mathbb{P}^{1}$-bundle $\mathbb{P}(E)$ over the Jacobian $J(X_{0})$, where $E$ is given by direct sum of degree zero line bundles on desingularization ${J}(X_{0})$. In [Chi19, Theorem 5.4], Chintapalli applied his result, Corollary 4.5 to desingularization of compactified Jacobians by setting $X=J(X_{0})$ with the direct sum $E$ of degree zero line bundles.
In the last section, §5, we provide an application of our result to desingularization of compactified Jacobians, generalizing [Chi19, Theorem 5.4].
@@
We work on the field of complex numbers $\mathbb{C}$. We use standard notation for abelian varieties. If $X$ is an abelian variety, we denote by $\hat{X}$ its dual abelian variety. If $X$ is an abelian variety, we denote by $\hat{X}$ its dual abelian variety. Denote by $\mathscr{P}$ the reduced Poincaré bundle on $X\times\hat{X}$. If we view $\hat{X}$ as the moduli spaces of topologically trivial line bundles on $X$, we have that $\mathscr{P}$ is the unique universal family such that $\mathscr{P}_{\{0\}\times\hat{X}}\simeq\mathcal{O}_{\hat{X}}$ and $\mathscr{P}_{X\times\{\mathcal{O}_{X}\}}\simeq\mathcal{O}_{{X}}$. Given an element $\alpha\in\hat{X}$, we let $P_{\alpha}$ be the corresponding line bundle in $\operatorname{Pic}^{0}(X)$, that is $P_{\alpha}=\mathscr{P}_{|X\times\{\alpha\}}$. Throughout this paper, we may use $\alpha$ to indicate $P_{\alpha}$. Given a integer $n$, we let $n_{X}:X\rightarrow X$ be the "multiplication by $n$" map. When $n$ is positive, we denote by $X[n]\coloneqq\operatorname{Ker}n_{X}$: the group-subscheme of $n$-torsion points of $X$. For every $x\in X$ we let $t_{x}:X\rightarrow X$ be the map defined by $a\mapsto a+x$. Given a polarization $l$ on $X$, we let $\varphi_{l}:X\rightarrow\hat{X}$ be the isogeny defined by $x\mapsto t_{x}^{*}L\otimes L^{-1}$, where $L$ is an ample line bundle with $c_{1}(L)=l$. It is a standard fact about abelian varieties that this does not depends from the choice of $L$. We will denote the kernel of $\varphi_{l}$ by $K(l)$. The identity element of $X$ is denoted simply by 0.
@@ 2.1. Syzygies of projective varieties
Let $Z$ be an algebraic variety over $\mathbb{C}$ and let $L$ be an ample invertible sheaf on $Z$. Let $S_{L}$ be the symmetric algebra over $H^{0}(Z,L)$. Then section ring associated to $L$,
@@
is a finitely generated graded $S_{L}$-module and as such it admits a minimal free resolution
(2.1) $F_{\bullet}=0\rightarrow\cdots\xrightarrow{f_{p+1}}F_{p}\xrightarrow{f_{p}}\cdots\xrightarrow{f_{2}}F_{1}\xrightarrow{f_{1}}F_{0}\xrightarrow{f_{0}}R_{L}\rightarrow 0$
with $F_{i}\simeq\bigoplus_{j}S_{L}(-a_{ij})$, $a_{ij}\in\mathbb{Z}_{>0}$. Following Green [Gre84a], we say that $L$ satisfies property $N_{p}$ for some nonnegative integer $p$, if, in the notations above,
$F_{0}=S_{L}$
@@ 2.2. $\mathbb{Q}$-twisted sheaves
$V^{i}\left(X,\mathcal{F}\left\langle\frac{a}{b}l\right\rangle,L\right)\coloneqq\{\alpha\in\hat{X}\>|\>h^{i}(X,b_{X}^{*}\mathcal{F}\otimes L^{ab}\otimes P_{\alpha})\neq 0\}.$
If we change the representative of $l$ we obtain a new locus that is a translate of $V^{i}\left(X,\mathcal{F}\langle\frac{a}{b}l\rangle,L\right)$ by an element in $\hat{X}$. In particular, the dimension of these loci does not depend from the choice of the line bundle $L$. If we change the representation of $\frac{a}{b}$, we have that still the dimension of these loci is unchanged. In fact, let $k$ be a positive integer. We have that
$(kb)_{X}^{*}\mathcal{F}\otimes L^{k^{2}ab}$ $\simeq k_{X}^{*}b_{X}^{*}\mathcal{F}\otimes\left(L^{ab}\right)^{k^{2}}$
$\simeq k_{X}^{*}\left(b_{X}^{*}\mathcal{F}\otimes L^{ab}\otimes\beta\right)$
for $\beta\in\hat{X}$ such that $(-1)_{X}^{*}L^{\frac{k^{2}-k}{2}}\otimes\beta^{k}\simeq L^{\frac{k^{2}-k}{2}}$. We deduce that $V^{i}\left(X,\mathcal{F}\left\langle\frac{ka}{kb}l\right\rangle,L\right)$ is the image of a translate of $V^{i}\left(X,\mathcal{F}\left\langle\frac{a}{b}l\right\rangle,L\right)$ through the isogeny $k_{X}^{*}$.
Thus we can introduce the following definition.
@@
Let $x$ a rational number. We say that $\mathcal{F}\left\langle xl\right\rangle$ is GV if $\operatorname{codim}V^{i}(X,\mathcal{F}\left\langle\frac{a}{b}l\right\rangle,L)\geq i$ for every $i\geq 1$, and for every representation $x=\frac{a}{b}$.
We say that $\mathcal{F}\left\langle xl\right\rangle$ is M-regular if $\operatorname{codim}V^{i}(X,\mathcal{F}\left\langle\frac{a}{b}l\right\rangle,L)\geq i+1$ for every $i\geq 1$, and for every representation $x=\frac{a}{b}$.
We say that a $\mathbb{Q}$-twisted sheaf $\mathcal{F}\left\langle\frac{a}{b}l\right\rangle$ satisfies I.T.(0) if its cohomological support loci are empty for every $i>0$and for every representation $x=\frac{a}{b}$.
@@ 2.3. Cohomological rank functions on abelian varieties
We see that the value of $h^{i}_{\mathcal{F},l}(x)$ depends only of $\mathcal{F}\langle xl\rangle$ ([JP20, Remark 2.2]). Furthermore, if a $\mathbb{Q}$-twisted sheaf $\mathcal{F}\langle xl\rangle$ satisfies I.T.(0) then we have that $h^{i}_{\mathcal{F},l}(x)=0$. In fact, let $x=\frac{a}{b}$ and chose $L$, such that $c_{1}(L)=l$. Then for all $\alpha\in\hat{X}$ we have that
$h^{i}(X,b_{X}^{*}\mathcal{F}\otimes L^{ab}\otimes\alpha)=0,$
when $i>0$. In particular $h^{i}(X,b_{X}^{*}\mathcal{F}\otimes L^{ab})_{\text{gen}}=0$ and $h^{i}_{\mathcal{F},l}(x)=0$. In addition to this, cohomological rank functions have been used to define invariants associated to polarizations of abelian varieties. Let $p\in X$ a closed point, we denote by $\mathcal{I}_{p}$ its associated sheaf of ideals. Given $l$ a polarization on $X$, we have that the quantity
@@
(2.6) $k(l)=\frac{\beta(l)}{1-\beta(l)}$
when $l$ is base-point free (that is when $\beta(l)<1$).
@@ 3. Ampleness of Tautological Bundles
Let $Z$ be a smooth projective complex variety. Let $V$ be a vector bundle of rank $\mathfrak{r}$ on $Z$. Let $\mathbb{P}_{Z}(V)$ be the associated projective bundle with the projection morphism $\pi:\mathbb{P}_{Z}(V)\rightarrow Z$. Let $\mathcal{O}_{\mathbb{P}(V)}(1)$ be the tautological line bundle of $\mathbb{P}_{Z}(V)$.
Definition 3.1.
A generalization of this notion is the concept of $k$-jet ampleness [BDRS99, §1.2].
@@
A vector bundle $V$ is 00-jet ample if and only if it is globally generated. In addition, according to [BDRS99, Proposition 4.2], if $V$ is a $1$-jet vector bundle on $Z$, then $V$ is very ample on $Z$.
Given a very ample vector bundle $V$ on $Z$, it make sense to give conditions ensuring property $N_{p}^{r}$ holds for the embedding induced by $\mathcal{O}_{\mathbb{P}(V)}(1)$. In [Par07] a cohomological criterion for property $N_{p}^{r}$ for the tautological bundle of a projective bundle $\mathbb{P}(V)\rightarrow Z$ is given. Before recalling the statement we need to fix some notation.
To this aim let $V$ be a very ample vector bundle on $Z$. In particular, we know that $V$ is globally generated and we have a short exact sequence
$0\rightarrow M_{V}\longrightarrow H^{0}(Z,V)\otimes\mathcal{O}_{Z}\longrightarrow V\rightarrow 0,$
where $M_{V}$ denotes the kernel of the evalutaion map $H^{0}(Z,V)\otimes\mathcal{O}_{Z}\rightarrow V$.
@@
This is essentially [Par07, Prop. 3.2]. Let $X=\mathbb{P}(V)$ and denote by $L$ the tautologiacal bundle on $X$, and let $\pi:X\rightarrow Z$ the morphism giving the projective bundle structure. It is well known that property $N_{p}^{r}$ is implied by the vanishing
$H^{1}\left(X,\bigwedge^{i}M_{L}\otimes V^{\otimes h}\right)=0$
for every $h\geq r+1$ and for every $11\leq i\leq p+$ 1. In the proof of [Par07, Prop. 3.2], it is shown how that this is implied by the vanishing
(3.1) $H^{1}\left(X,\bigwedge^{l}\pi^{*}M_{V}\otimes\Omega^{m}_{X/Y}\otimes L^{\otimes h}\right)=0,$
for every $l+m=i$ and for $h\geq r+1$. As a consequence of projection formula plus an easy cohomological argument, we have that (3.1) is implied by the vanishing
(3.2) $H^{k}\left(Z,\bigwedge^{l}M_{V}\otimes\bigwedge^{m+1+k}V\otimes S^{m+h-k}V\right)=0,$
for every $k\geq 1$, $l+m=i$, and for $h\geq r+1$. Now, $\bigwedge^{m+1+k}V\otimes S^{m+j-k}V$ is a summand of $V^{\otimes m+j}$, therefore we see that (3.2) is implied
@@
for every $l+m=i$ and every $h\geq r+1$. ∎